OpenStudy (goformit100):

A and B two candidates seeking addmission in IIT. The probability that A is selected is 0.5 and the probability that both A and B are selected is atmost 0.3. Is it possible that the probability of B getting selected in 0.9 ?

OpenStudy (goformit100):

@oldrin.bataku

OpenStudy (anonymous):

The events that A gets selected or B gets selected are independent. Hence:\[\bf P(A \ and \ B) = P(A)P(B)=0.5P(B)\]Now the probability P(B) is atmost 0.3 hence:\[\bf 0<0.5P(B) \le 0.3 \implies 0 < P(B) \le 0.6\]Hence it is not possible that the probability of B getting selected is 0.9. @goformit100

OpenStudy (anonymous):

Sorry I meant to say: "Now the probability of BOTH B and A getting selected is atmost 0.3 hence:" Replace the second phrase with this.

OpenStudy (goformit100):

Hey i got the answer Yu huuu

OpenStudy (anonymous):

the question asks for "is it possible that B has a 0.9 probability of getting selected. I don't see how that answers the question? lol..

OpenStudy (goformit100):

Any ways Thanks bro.

OpenStudy (anonymous):

yw